Gold crystallizes in ccp structure.The total no. of voids present in 197g of gold will be{Au=197}???

Grade:12th pass

4 Answers

Suraj Prasad IIT Patna
7 years ago

Number of atoms of Au = 197/197×N(A). In N close-packed spheres there are 2N tetrahedral interstices and N octahedral interstices for a total of 3N interstices. The answer is ∴ 3N(A)
Sayan Ghosh
6 years ago
For 4 Au atoms in unit cell we get 12 voids .so for 1 atom we will get 12/4 voids.hence for Na no of atoms we will get 12/4 ×Na no of voids.and the answer is 3 ×Na no of voids
Oshin
6 years ago
w/m of Au = 197 / 197 = 1 No. of unit cells = 4 (in ccp)Voids in octahedral and tetradral sites is = 2n + n = 3n •Total voids = 4× 3 = 12 Hence, for 1 atom = 12 / 4 =3 Na
Kushagra Madhukar
3 years ago
Dear student,
Please find the attached solution to your problem.
 
Number of atoms of Au = 197/197×NA.
In 1 close-packed spheres there are 2 tetrahedral interstices and 1 octahedral interstices.
Hence total number of voids = 2NA + NA = 3NA
∴ The answer is 3NA
 
Hope it helps.
Thanks and regards,
Kushagra